Assume that you are considering the purchase of a 10-year bond with a coupon rate of 10%

 14

 Assume that you are considering the purchase of a 10-year bond with a coupon rate of 10%. The bond has a par value of $1,000 and makes semiannual interest payments.

 The bond sels for $894.50, and can be called in 4 years for S1100. What is the yield to call of this bond?

  •  15.54%

  •  11.49%

  •  12.85%

  •  11.83%

  •  13.77%


 15

 The expected return on some company's stock is 14%. The stock's dividend is expected to grow at a constant rate of 8%, and it currently sells for S50 a share. Which of the following statements is CORRECT

  •  The stock's dividend yield is 8%.

  •  The stock's dividend yield is 6%.

  •  The stock price is expected to be $50 a share one year from now.
Answer #1

14

Calculator
Inputs:
PV         (894.50)
PMT             50.00
FV       1,100.00
N                     8
Output:
I/Y = Period rate 7.77%
APR= I/Y *2 15.54%

Answer is:

15.54%

15

Dividend next year = 50 * (14%-8%) = 3

Dividend yield = 3/50 = 6%

Answer is:

The stocks dividend yield is 6%

Please rate.